A random sample of 92 observations produced a mean x=25.4 and a standard deviation =2.8.

a. Find a​ 95% confidence interval for μ.

b. Find a​ 90% confidence interval for μ.

c. Find a​ 99% confidence interval for μ.

a. The​ 95% confidence interval is (____.______)

​(Use integers or decimals for any numbers in the expression. Round to two decimal places as​ needed

b. The​ 90% confidence interval is (____.______)

​ (Use integers or decimals for any numbers in the expression. Round to two decimal places as​ needed

c.The​ 99% confidence interval is (____.______)

​ (Use integers or decimals for any numbers in the expression. Round to two decimal places as​ needed
Answer #1


Solution :

Given that,

\bar x =25.4

s =2.8.

n = 92

Degrees of freedom = df = n - 1 = 92 - 1 = 91

a) At 95% confidence level the z is ,

\alpha = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

t\alpha /2,df = t0.025,91 =1.986

Margin of error = E = t\alpha/2,df * (s /\sqrtn )

= 1.986 * (2.8. / \sqrt 92)

= 0.58

Margin of error = 0.58

The 95% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

25.4 - 0.58 < \mu < 25.4 + 0.58

24.82 < \mu < 25.98

(24.82, 25.98 )

b ) At 90% confidence level the z is ,

\alpha = 1 - 90% = 1 - 0.90 = 0.10

\alpha / 2 = 0.10 / 2 = 0.05

t\alpha /2,df = t0.05,91 =1.662

Margin of error = E = t\alpha/2,df * (s /\sqrtn )

= 1.662 * (2.8. / \sqrt 92)

= 0.49

Margin of error = 0.49

The 90% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

25.4 - 0.49 < \mu < 25.4 + 0.49

24.82 < \mu < 25.89

(24.91, 25.89 )

c ) At 99% confidence level the z is ,

\alpha = 1 - 99% = 1 - 0.99 = 0.01

\alpha / 2 = 0.01 / 2 = 0.005

t\alpha /2,df = t0.025,91 =1.986

Margin of error = E = t\alpha/2,df * (s /\sqrtn )

= 1.986 * (2.8. / \sqrt 92)

= 0.77

Margin of error = 0.77

The 99% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

25.4 - 0.77< \mu < 25.4 + 0.77

24.63 < \mu < 26.17

(24.63, 26.17 )